OpenStudy (anonymous):

Sophia has 8 books in her locker. All the books are either school books or personal books. She has three times as many school books than personal books. How many school books does Sophia have in her locker?

OpenStudy (anonymous):

Total books = School + Personal Total = 8 Personal = 3*S School = S Now you try to finish it, what do you get?

OpenStudy (campbell_st):

x = personal books.......3x = school books sum of the books is 8 then x + 3x = 8 4x = 8 x = 2 then there are 2 personal books and 6 school books

OpenStudy (anonymous):

Thanks !